OpenStudy (anonymous):

I bought 32 cans. .80c for tomatoes .40c for corn. I spent $18. How much corn did I buy? 19?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@Hero

OpenStudy (amistre64):

c + t = 32 .8t + .4c =18 what should we do?

OpenStudy (amistre64):

we want to know c, let c=n, t = 32-n

OpenStudy (amistre64):

.8(32-n) + .4n =18 when we known we know c.orn